Shift report 20010911 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Sep 11 2001 - 11:40:47 EDT

  • Next message: Flemming Videbaek: "RHIC and BNL status"

    ------------------------------------------------------------
    Date: Tue Sep 11 11:40:46 2001, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#8 pages 112 - 113
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from 8:44 to 9:21, about 40min.
    
    2. Run 4696 : Data run 138k trigger 1
       Run 4702 : Camac Pedestal Run
       Run 4703 : TPC Pedestal Run
    
    3. No detector problems.
       During the data-taking for Run 4696, T3 tripped once.
       The currents of T3-T5 as follows;
       
       T3 : Anode 1 (110.72)
            Anode 2 ( 88.96)
            Anode 3 ( 96.52)
       T4 : Anode 1 ( 26.57)        
            Anode 2 ( 19.47)
            Anode 3 ( 17.00)
       T5 : Anode 1 (118.14)
            Anode 2 ( 56.10)
            Anode 3 ( 63.30)
    
       T3 Field(1-3) Actual values are fluctuated. (up and down)
    
       RHIC message:C-A Department facilities will be shut down 
                    for the next 24 hrs.
    
       Magnet D1-D5 : standby state     
     
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Sep 11 2001 - 11:41:21 EDT